OpenStudy (anonymous):

Hello, can somebody help me? I currently need to solve a related rate problem and I want to know the general way to do it. How do I solve for the rate of volume increasing/decreasing when the height increases at a certain rate and the length decreases at a certain rate?

OpenStudy (anonymous):

OOOOhhhh lemmer seee

OpenStudy (anonymous):

look and see if this kinda helps dnichols30582.edublogs.org/files/2014/06/3.11-Solutions-2i0f245.pdf

OpenStudy (anonymous):

www2.bc.cc.ca.us/resperic/math6a/lectures/ch4/4/HW4.4/HW4.4sols.htm

OpenStudy (inkyvoyd):

you might want to be specific. I can try to answer it generally but that might confuse you more.

OpenStudy (inkyvoyd):

oh wait you asked for generalization sec

OpenStudy (inkyvoyd):

How do I solve for the rate of volume increasing/decreasing when the height increases at a certain rate and the length decreases at a certain rate? you want the rate of volume change. Or, dV/dt V is a function of multiple things... V(l,h)

OpenStudy (kainui):

I'm not watching but maybe I'll comment in an hour from now when I'm done browsing dank memes

OpenStudy (anonymous):

Bless @Kainui

OpenStudy (anonymous):

In my case, the height is 4 and increases by 2 cm/min and the height is 15 cm and decreases at 3 cm/min. I just wanna learn the general way to do it so get a better conceptual understanding, you know?

OpenStudy (inkyvoyd):

what comes next might be considered multivariate calculus (I believe it's the multivariate chain rule) but let's see generally speaking l is a function of h, or h is a function of l (say in the case of a cone, there is a Pythagorean relationship etc) so we can rewrite as V(l,h(l)).... we want \(\Huge \frac{dV}{dt}=\frac{\partial V}{\partial l}*\frac{dl}{dt}+\frac{\partial V}{\partial h}*\frac{dh}{dt}\) or \(\Huge \frac{dV}{dt}=\frac{\partial V}{\partial l}*\frac{dl}{dt}+\frac{\partial V}{\partial h}*\frac{dh}{dl}*\frac{dl}{dt}\)

OpenStudy (inkyvoyd):

note the multivariate chain rule states, for two variables (extendable up to n), that if f is a function of x and y, and x and y are functions of s, then \(\Huge \frac{df}{ds}=\frac{\partial f}{\partial x}*\frac{dx}{dt}+\frac{\partial f}{\partial y}*\frac{dy}{dt}\) partial symbols which look like backward sixes just mean that you derive with respect to that variable in particular whilst holding other variables to be constant, for instance; \(\Huge \frac{\partial}{\partial x}(5x^3y^2za)=15x^2y^2za\)

OpenStudy (inkyvoyd):

Correction: those t's should be s's sorry

OpenStudy (inkyvoyd):

Let me know if you need clarifications.

OpenStudy (anonymous):

I'm trying to apply this to the math question I'm currently on. Thank you.

OpenStudy (inkyvoyd):

Note that if h is not a function of l than the first form \(\Huge \frac{dV}{dt}=\frac{\partial V}{\partial l}*\frac{dl}{dt}+\frac{\partial V}{\partial h}*\frac{dh}{dt}\) is applicatble and not the second as dh/dl is non sequitur